2. (60 points total) Please show all work Suppose the labor supply for a firm is given by: and REAL Labor Demand is given by. ND 120 2N a)(5 points) What is the market clearing real wage and level of employment? b)lS points) Suppose the government imposed a minimum real wage of 30. How many workers are willing to work at a real wage of 30?
Answer #1

a) The market clearing real wage equates the supply and demand of labor.

Let's set w= MPN = 120-2N

Solving for N gives

2N= 120 - w

Therefore N = 60 - 0.5w

Equating labor demand to supply

60 - 0.5w = 10 + 2w

2.5w = 50

Hence w = 20

Market clearing real wage is 20.

(b) Minimum wage of 30 , it is above the market wage.

At w = 30, labor demand is 60 - 0.5w

that is 60 - 0.5(30) = 45

and labor supply is 10 + 2w that is

10 + 2(30) = 70

So unemployment is 70 - 45 = 25 workers